### v1

# Drift time calibration of the ultra-Low material budget GEM-based TPC for MIXE

Abstract
Muon-Induced X-ray Emission (MIXE) is a non-destructive analytical technique that leverages negative muons to probe elemental and isotopic compositions by detecting characteristic muonic X-rays emitted during atomic cascades and gamma rays from nuclear capture processes. By controlling the muon beam momentum, MIXE enables depth-resolved analysis, spanning microns to centimeters, making it ideal for studying compositional variations in fragile, valuable, or operando samples. To enhance its capabilities, we integrated a twin Time Projection Chamber (TPC) tracker with Gas Electron Multiplier (GEM) amplification stages, allowing precise measurement of muon trajectories. A custom-built fiber detector with scintillating fibers and a Silicon Photomultiplier (SiPM) provided sub-percent-level accuracy in drift velocity calibration, essential for accurate spatial reconstruction. This advanced setup correlates muon stopping points with X-ray emissions, paving the way towards element-sensitive imaging and establishing MIXE as a unique tool for high-resolution, depth-specific elemental analysis across diverse scientific applications.
1 Introduction
Muon-Induced X-ray Emission (MIXE) [1, 2, 3, 4] is a powerful technique that leverages exotic atom formation by irradiating target samples with a well-defined beam of negative muons. As muons stop within the material, they are captured by target atoms, forming excited atomic states and cascading to the ground state on a picosecond timescale [5, 6]. During this process, X-rays are emitted at characteristic energies specific to the atomic species, enabling precise identification of both elemental and isotopic compositions. In addition to atomic cascades, muons can also undergo nuclear capture, a process that becomes more probable in heavier elements [7]. During nuclear capture, the muon catalyzes the conversion of a proton into a neutron, resulting in a transmuted nucleus. These nuclei de-excite by emitting a combination of nucleons and gamma rays. The combination of muonic X-rays from atomic cascades and gamma rays from nuclear interactions provides multiple pathways for elemental identification. High-purity germanium (HPGe) detectors are used to measure these emissions with excellent energy resolution and efficiency over a broad spectrum, ensuring accurate and detailed analysis.
One significant benefit of MIXE over similar techniques, like X-Ray fluorescence (XRF) [8], is its depth-sensitivity. Muons, as charged particles, are implanted with well defined stopping points, spanning from microns to centimeters, dependent on their initial momentum. By adjusting the muon beam’s momentum, MIXE enables depth-specific elemental analysis, aiding the examination of compositional changes from the surface to deeper layers of a material. Being fundamentally non-destructive, MIXE is highly suitable for assessing valuable or rare samples. Consequently, MIXE is an excellent method for studying precious samples, such as cultural heritage artifacts [9], or performing operando studies, such as charging kinetics in batteries [10]. Building on MIXE’s core advantages, we aim to extend the technique by merging the exact muon stopping point data with the corresponding X-ray information, facilitating a new non-destructive, element-sensitive tomographic method.
To facilitate this development, we adapted a twin Time Projection Chamber (TPC) tracker equipped with Gas Electron Multiplier (GEM) amplification stages, originally developed for heavy-ion tracking [11]. During the beamtime at the Paul Scherrer Institute (PSI), the TPC was positioned upstream of the target to capture the precise muon trajectory. The X-position of the muon is determined by the projection of charge clusters onto the strips, while the Y-position is derived from the drift time of the electrons. Thus, an accurate calibration of the drift velocity in the gas is crucial for ensuring the reliability of the position reconstruction. Since this calibration needs to be conducted in situ, we developed an independent compact fiber detector and successfully employed it to calibrate the drift velocity with high accuracy, achieving better than percent-level precision.
2 Drift time calibration
The calibration detector (see Figure 1) consists of a precision 3D-printed (35 m resolution) mounting structure, housing three scintillating optical fibers of approximately 50 mm length, with 1x1 mm² cross-section, spaced exactly 4 mm apart. The active are of the fibers are shielded from light by two thin (10 µm) titanium foils, allowing low energy muons to enter with minimal energy loss. The rest of the casing is sufficiently thick that muons are absorbed in the material, before reaching the fibers. The signal is read out by a high-speed Silicon Photomultiplier (SiPM), which is mounted on an optimized readout board (onsemi MICROFJ-SMA-60035-GEVB), using the fast signal output, which is amplified by two Minicircuits ZFL-1000LN+ amplifiers in series.
The optical fiber detector was placed in front of the exit window of the twin GEM-TPC to simultaneously detect incoming muons in multiple systems: the entrance detector [12], the tracker, and one of the fibers in the optical fiber detector. A stringent cut on the total drift time sum in both GEM-TPCs allows selecting parallel tracks for which precise tracking of muon trajectories through the system becomes possible. During data analysis, events recorded by the GEM-TPC that correspond to hits in one of the three fibers can then be identified.
This correlation enables the generation of a drift time distribution with three distinct peaks, each representing the interaction of a muon with one of the fibers. The peaks are then fitted with a sum of three coupled Gaussian functions:
where the model assumes that the Gaussian peaks are symmetrically distributed around the central time , with offset timing . The parameters and represent the amplitudes, and and represent the widths of the respective distributions. Since the separation between directly corresponds to the offset timing, it is then simply given by
3 Results
During the PSI beam experiments, the chamber was evaluated using two gas mixture: Ar/CO2 (75:25), and He/CO2 (90:10). High voltages were applied to the GEM-TPCs, with 300 V/cm for the drift regions and 3.01 kV (top) and 3 kV (bottom) applied to the triple GEM stacks for Ar/CO2, and 325 V/cm for the drift regions and 2.16 kV (top) and 2.15 kV (bottom) applied to the triple GEM stacks for He/CO2.
Utilizing a muon beam with moment of 50 (He/CO2) to 60 (Ar/CO2) MeV/c, the drift times for both the top and bottom GEM-TPCs were evaluated, as depicted in Figure 2 and Figure 3.Using these measurements and equation 2.2, the drift velocities for Ar/CO2 gas mixture were calculated to be and , respectively. For the He/CO2 gas mixture, the drift velocities were similarly calculated to be and .
This in-situ calibration method provides an efficient and precise approach to determine the tracker’s drift velocity, eliminating the need for additional large-scale detectors. The measurements are consistent with results obtained with high energy muons using a reference tracker system [13]. This capability will be crucial to ensure reliable operation, while being able to change the parameters of the GEM-TPCs to cover the range of muon momenta required for depth-selective studies.
Acknowledgments
This work is based on experiments performed in September 2023 and June 2024 at the PiE1 beamline of the Swiss Muon Source SµS (proposals 20222993 and 20231296), Paul Scherrer Institute, Villigen, Switzerland. The authors gratefully acknowledge the Laboratory for Particle Physics (LTP) and the Laboratory for Muon Spin Spectroscopy (LMU) at PSI for their generous support. We also extend our gratitude to members of the CERN GDD team and of RD51/DRD1 test beam working group for their invaluable contributions. Many thanks to C. Kaya, J. Kunkel, B. Voss, H. Risch, C. Schmidt, R. Turpeinen, J. Heino, S. Rinta-Antila, R. de Oliveira, and the CERN Micro Pattern Technology (MPT) Workshop for their support during the design and assembly of the TPC. This work was supported by the SNSF bilateral program grant 214052, SNSF Sinergia grant 193691 and the SDSC collaborative project SAMURAI (C22-12L).